# CostBench: Evaluating Multi-Turn Cost-Optimal Planning and Adaptation in Dynamic Environments for LLM Tool-Use Agents

### v1

@@ 5.1 Static Setting
In the unblocking setting, the model’s task is to enumerate all possible paths and select the optimal one. We define coverage as the proportion of paths explicitly planned, and hypothesize that higher coverage leads to better performance. To identify distinct paths in each model’s output, we use GPT-4o-mini OpenAI (2024a) (prompt in Figure 11). As shown in Table 4, performance follows GPT-5 >> Gemini-2.5-pro >> open-source Qwen models, which aligns with the coverage trend in Figure 4. Performance also declines as task sequences grow longer (Figures 2), explained by the corresponding drop in coverage. Importantly, even when given abstract examples that enumerate all paths (Figure 15), models often fail to generalize these planning strategies to real scenarios.
@@ 5.2 Dynamic Setting
Multiple blockings severely undermines model robustness.
We further analyze the results under multiple blocking events (Figure 6), focusing on how repeated disruptions impact model robustness and cost-optimal planning. We find that increasing the frequency of both ban tool and cost change blocks leads to a substantial drop in EMR across all models. For instance, the EMR of Gemini-2.5-Pro, which initially achieved around 45% accuracy, drops to nearly 0.1 after three consecutive cost-change events, while Qwen3-14B completely fails (EMR = 0) after only two such events. This indicates that even state-of-the-art commercial and open-source models struggle to maintain reasoning consistency and cost-awareness under dynamically perturbed environments, highlighting their limited robustness and adaptability to repeated disruptions.
@@ 6 Conclusion
While CostBench provides a principled and scalable framework for evaluating cost-aware planning, several limitations remain. First, our current tasks are restricted to the travel-planning domain, which, though diverse, may not fully capture the breadth of real-world cost trade-offs. Nevertheless, CostBench can be easily extended to other domains by adding configuration files. Second, our simulation abstracts away true API latency, resource consumption, and stochastic failures, which may further influence agent behavior in deployment. Third, the blocking events we design, while representative, are still manually parameterized; future work could explore learning-based or user-driven dynamics to better approximate natural environmental shifts. Finally, incorporating multimodal agents capable of processing and reasoning over diverse data types such as text and images represents another critical research direction Su et al. (2025b).
@@
Our code will be released under the MIT license to allow unrestricted research use. The CostBench will be distributed under a Creative Commons (CC) license, providing free access for the academic community. We take full responsibility for any potential rights violations or licensing issues, and all resources comply with their respective terms of use while supporting research purposes.
@@ Appendix A Comparison Traits Details
  • Tool Use: The benchmark incorporates or necessitates the use of external tools. This trait is particularly important, as tool use has emerged as a defining capability of modern LLM agents—central to their ability to interact with and act upon the external world Wang et al. (2024b); Qin et al. (2024b). Evaluating tool use is thus indispensable for understanding an agent’s true reasoning and decision-making competence.
  • Dynamic State: The environment introduces runtime-controllable obstacles that may block the agent and enforce replanning. This design is important because traditional benchmarks often suffer from data leakage risks Guo et al. (2025); Wu et al. (2025b). In contrast, a dynamic environment can mitigate such risks by introducing runtime factors during evaluation, ensuring a fairer and more reliable assessment.
  • Adjustable Difficulty: The difficulty level can be systematically adjusted, for example by scaling task parameters. A naturally adjustable difficulty hierarchy allows systematic evaluation of model adaptability under varying levels of challenge Qian et al. (2025b); Zong et al. (2025b); Huang et al. (2025a), facilitating more fine-grained analysis.
@@
  • Customized Details: The environment offers fine-grained configurability of task details. Enabling this feature makes the benchmark more adaptable to other domains and facilitates broader applicability across diverse evaluation settings Lu et al. (2025); Wang et al. (2024b). Scalability: The benchmark can be automatically expanded in size without additional human annotation and easily generalized to other tasks. This property is crucial, as scalability and expandability are essential for large-scale evaluation Qian et al. (2025b); Yao et al. (2025); Zong et al. (2025a); Wang et al. (2025e).
@@ Appendix B Experiment Details
    In this section, we provide a comprehensive overview of the experimental setup and evaluation methodology. We begin by detailing the construction of tools, including their names and descriptions. Next, we explain the simulation-based derivation of greedy and ground-truth trajectories using a tool graph framework, covering both non-blocking and blocking scenarios. We then present the evaluation metrics, including cost gaps, path dissimilarities, user intent alignment, and tool call validity, along with their formulas and boundary conditions. Finally, we describe the mechanisms for implementing blocking events, prove the solvability of blocking scenarios, and outline the randomized yet reproducible environment generation process to ensure fairness and reproducibility in evaluations. An illustration of our environment is shown in Figure 1, while the primary hyperparameter configurations are presented in Table 5.
@@ B.1 Tool Construction Details
    To better evaluate the model’s genuine planning ability, we mitigate potential overfitting to existing data Wu et al. (2025b); Liu et al. (2025). Since biases acquired during training may distort the model’s assessment of tool cost-effectiveness, we construct a new set of tools for evaluation. In this section, we mainly introduce two aspects of the tool schema: name construction and description construction.
@@ B.2 Greedy and Ground-truth Calculation
    This construction allows us to simulate greedy and ground-truth trajectories over the tool graph without actually executing tools. While such stimulation offers a convenient and controllable proxy, it fundamentally differs from real tool use: being algorithm-driven, it bypasses the core challenge of user-intent understanding. Consequently, although useful for isolating certain planning aspects, it fails to deal with the vagueness, ambiguity, and negotiation with human preferences inherent in real scenarios. Our benchmark therefore becomes essential: cost-aware planning cannot be addressed by traditional algorithmic stimulation, but calls for evaluation of large language models, which uniquely possess the capacity to reason over ambiguity and align with human intent Yao et al. (2025).
@@ B.3.1 Detailed Implementation and Formulas
    User Intent Metric.
    We report User Intent Hit Ratio as the proportion of samples where the agent’s predicted final candidate exactly matches the ground-truth answer. This captures agent’s ability to understand and intepret human intent in natural language.
    Tool Call Metrics.
@@ B.3.2 Boundary Conditions
    For path-related metrics, we focus on evaluating genuine cost-optimal planning ability. Trajectories that fail to reach the goal state are excluded from metric computation to ensure that cost awareness is measured only on successful completions. To avoid being influenced by agents’ tool calling ability and reflect true cost-optimal planning capability, we also exclude failed tool calls. We thereby disentagle agents’ tool calling ability and cost-optimal planning ability.
    User Intent Metrics.
    For User Intent Hit Ratio, we exclude all samples in which the agent fails to reach the goal state, since such predictions amount to mere guesses rather than answers derived from a complete tool-call trajectory.
@@ B.6.1 Deterministic Cost Assignment
    Here $a$ denotes an atomic tool, $T$ a composite tool with $k$ atomic components, and $h(\cdot)$ the hash-based seed derivation function. The $\sqrt{k}$ scaling in equation (2) reflects empirical observations that execution variance grows sublinearly with pipeline length.
@@ B.6.2 Deterministic Blocking Parameters
    Repeated Calls $\downarrow$ Extra Calls $\downarrow$ Wrong Parameters $\downarrow$ Unaccessible Calls $\downarrow$
@@
    Among the four types of blocking events, all trigger times are determined based on the optimal path. Therefore, given identical cost assignments, the trigger times remain fixed. The blocking parameters are consistent across all blocking types except for ban tool, as they are uniquely controlled by the seed $S_{q,i}$. In contrast, the ban tool event is model-dependent—it is triggered at a specified time, and whichever tool the model invokes at that timestep will result in a failure.
    Although this design means the environment is not perfectly identical across models, we argue that it does not compromise fairness. If a model follows the correct trajectory, both the ground truth and the agent are equally affected after the ban event, requiring replanning and thus preserving fairness in evaluation. Conversely, if a model already deviates from the optimal path, the model-dependent blocking has no impact, as the datapoint’s EM score would already be zero. Consequently, the EMR metric remains unaffected. Moreover, as shown in Table 4, EMR and ANED/AED exhibit opposite trends, suggesting that the latter two effectively capture performance differences as well.
    Thus, this hierarchical seeding ensures that: (i) given $(S,q)$ and the blocking type, the entire blocking sequence and all associated parameters are fully determined; (ii) different queries exhibit different blocking patterns even with the same $S$; and (iii) experimental results are perfectly reproducible, facilitating controlled comparisons across models and further analysis. At the same time, users can increase the benchmark’s diversity by adjusting random seeds, allowing robustness evaluation under varied conditions and mitigating potential data contamination.
    By controlling a randomized yet reproducible environment, we ensure fair comparison across different runs and model configurations.
@@ Appendix C Additional Experiment Results
    C.1 The Effect of Random Seeds
    To ensure the robustness of our conclusions regarding the cost assignment, which is controlled by a random seed, we conducted three additional experiments using seeds 1000, 2000, and 3000 to evaluate model performance. As illustrated in Figure 8, the LLMs exhibit no significant performance variation across different seeds, with differences in ANED and EMR remaining within 5%, confirming low seed sensitivity. However, Figure 3 reveals that LLMs are highly sensitive to variations in cost assignment. We attribute the observed stability across seeds to the diversity and size of our test set, which mitigates the impact of seed-dependent cost assignments and ensures robust evaluation.

### v2

@@ 5.1 Static Setting
In the unblocking setting, the model’s task is to enumerate all possible paths and select the optimal one. We define coverage as the proportion of paths explicitly planned, and hypothesize that higher coverage leads to better performance. We employ GPT-4o-mini OpenAI (2024a) to extract distinct paths from each model’s output (prompt in Figure 12, human check in Appendix F). As shown in Table 3, performance follows GPT-5 >> Gemini-2.5-pro >> open-source Qwen models, which aligns with the coverage trend in Figure 4. Performance also declines as task sequences grow longer (Figure 2), explained by the corresponding drop in coverage. Importantly, even when given abstract examples that enumerate all paths (Figure 16), models often fail to generalize these planning strategies to real scenarios.
@@ 5.2 Dynamic Setting
Multiple blockings severely undermine model robustness.
We further analyze the results under multiple blocking events (Figure 6), focusing on how repeated disruptions impact model robustness and cost-optimal planning. We find that increasing the frequency of both ban tool and cost change blocks leads to a substantial drop in EMR across all models. For instance, the EMR of Gemini-2.5-Pro, which initially achieved around 45% accuracy, drops to nearly 0.1 after three consecutive cost-change events, while Qwen3-14B completely fails (EMR = 0) after only two such events. This indicates that even state-of-the-art commercial and open-source models struggle to maintain cost-awareness under complex dynamic environments, highlighting their limited robustness and adaptability to repeated disruptions.
@@ 6 Conclusion
While CostBench provides a principled and scalable framework for evaluating cost-aware planning, several limitations remain. First, our current tasks are restricted to the travel-planning domain, which, though diverse, may not fully capture the breadth of real-world cost trade-offs. Nevertheless, CostBench can be easily extended to other domains by adding configuration files. Second, our simulation abstracts away true API latency, resource consumption, and stochastic failures, which may further influence agent behavior in real-world deployment. Third, the blocking events we design, while representative, are still manually parameterized; future work could explore learning-based or user-driven dynamics to better approximate natural environmental shifts.
@@
Our code will be released under the MIT license to allow unrestricted research use. The CostBench will be distributed under a Creative Commons (CC) license, providing free access for the academic community. Our use of existing models and tools is strictly consistent with their original licenses and intended research purposes. We take full responsibility for any potential rights violations or licensing issues, and all resources comply with their respective terms of use while supporting research purposes.
@@ Appendix A Comparison Traits Details
  • Tool Use: The benchmark incorporates or necessitates the use of external tools. This trait is particularly important, as tool use has emerged as a defining capability of modern LLM agents—central to their ability to interact with and act upon the external world Lu et al. (2025); Wang et al. (2024b); Qin et al. (2024b); Huang et al. (2024). Evaluating tool use is thus indispensable for understanding an agent’s true reasoning and decision-making competence.
  • Dynamic State: The environment introduces runtime-controllable obstacles that may block the agent and enforce replanning. This design is important because traditional benchmarks often suffer from data leakage risks Guo et al. (2025); Wu et al. (2025b); Cheng et al. (2025); Singh et al. (2024). In contrast, a dynamic environment can mitigate such risks by introducing runtime factors during evaluation, ensuring a fairer and more reliable assessment.
  • Adjustable Difficulty: The difficulty level can be systematically adjusted, for example by scaling task parameters. A naturally adjustable difficulty hierarchy allows systematic evaluation of model adaptability under varying levels of challenge Qian et al. (2025c); Zong et al. (2025); Huang et al. (2025a), facilitating more fine-grained analysis.
@@
Tools Value
Task Sequence NN
Total Atomic Tools 6NN
Total Composite Tools 6N(N1)/2{N*(N-1)/2}
Total Tools 6N(N+1)/2{N*(N+1)/2}
Table 4: Main statistics of tools used in CostBench. The variable NN denotes an adjustable task sequence, and the total numbers of atomic and composite tools vary accordingly, as indicated (we use $N=5$ in most of our experiments).
@@
  • Customized Details: The environment offers fine-grained configurability of details. Enabling this feature makes the benchmark adaptable to other domains and facilitates broader applicability across diverse evaluation settings Lu et al. (2025); Wang et al. (2024b).
  • Scalability: The benchmark can be automatically expanded in size without additional human annotation and easily generalized to other tasks. This property is crucial, as scalability and expandability are essential for large-scale evaluation Qian et al. (2025c); Yao et al. (2025).
@@ Appendix B Experiment Details
    In this section, we provide a comprehensive overview of the experimental setup and evaluation methodology. We begin by detailing the construction of tools, including their names and descriptions. Next, we explain the simulation-based derivation of greedy and ground-truth trajectories using a tool graph framework, covering both non-blocking and blocking scenarios. We then present the evaluation metrics, including cost gaps, path dissimilarities, task completion rates, and tool call invalidity, along with their formulas and boundary conditions. Finally, we describe the mechanisms for implementing blocking events, prove the solvability of blocking scenarios, and outline the randomized yet reproducible environment generation process to ensure fairness and reproducibility in evaluations. An illustration of our environment is shown in Figure 1, while the primary hyperparameter configurations are presented in Table 5.
@@ B.1 Tool Construction Details
    To better evaluate the model’s genuine planning ability, we mitigate potential overfitting to existing data Wu et al. (2025b); Liu et al. (2025a). Since biases acquired during training may distort the model’s assessment of tool cost-effectiveness, we construct a new set of tools for evaluation. In this section, we mainly introduce two aspects of the tool schema: name construction and description construction.
@@
    Model Average Tool Call Numbers
    Qwen3-8B 2.18
    Qwen3-14B 3.42
    Qwen3-32B 2.84
    Llama-3.1-8B-Instruct 1.99
    GLM-4.5 2.86
    Deepseek-V3.1 3.11
    Gemini-2.5-Pro 2.75
    Claude-Sonnet-4 2.62
    GPT-4o 3.36
    GPT-5 2.69
    Table 6: The statistics of average tool call counts for all evaluated models in the static evaluation (in Table 3) include only valid tool calls; all invalid calls are excluded from the computation.
@@ B.2 Greedy and Ground-truth Calculation
    This construction allows us to simulate greedy and ground-truth trajectories over the tool graph without actually executing tools. While such stimulation offers a convenient and controllable proxy, it fundamentally differs from real tool use: being algorithm-driven, it bypasses the core challenge of user-intent understanding. Consequently, although useful for isolating certain planning aspects, it fails to deal with the vagueness, ambiguity, and negotiation with human preferences inherent in real scenarios. Our benchmark therefore becomes essential: cost-aware planning cannot be addressed by traditional algorithmic simulation, but calls for evaluation of large language models, which uniquely possess the capacity to reason over ambiguity and align with human intent Yao et al. (2025).
@@
    Note that while our randomized cost assignment strategy naturally minimizes the likelihood of identical path costs, ties may still theoretically occur. To ensure a unique ground truth, we implement a tie-breaking rule: among paths with equivalent total costs, the trajectory with fewer tool calls is selected as the optimal solution. This criterion is strictly enforced in our evaluation and is also explicitly communicated to the models via the inference prompt (see Figure 15).
    Figure 8: An illustrative example of metric calculation. The diagram contrasts the ground-truth trajectory ($\tau_{gt}$, indicated in green) with the agent’s trajectory ($\tau_{agent}$, indicated in red). The numbers on the edges represent the specific cost of each tool invocation. We use this scenario to demonstrate how Cost Gap, path-related metrics (AED, ANED, EMR), and task completion metrics are computed.
@@ B.3.1 Detailed Implementation and Formulas
    Another important aspect is that, by design, the environment assigns each composite tool a cost equal to the sum of its component atomic tools plus zero-mean Gaussian noise. As a result, the expected cost of any tool-call sequence reaching the goal state is identical, regardless of its length. Therefore, the cost gap cannot be reduced to a proxy for the number of tool calls; instead, it captures whether the model adopts cost-optimal invocation patterns under the given cost structure.
@@
    Task Completion Metric.
    We report Task Completion Ratio as the proportion of samples where the agent’s predicted final candidate exactly matches the ground-truth answer. This mainly captures the agent’s ability to understand and interpret human intent expressed in natural language, as well as to invoke the necessary tools to obtain the final answer.
    Tool Call Metric.
@@ B.3.2 Boundary Conditions
    For path-related metrics, we focus on evaluating genuine cost-optimal planning ability. Trajectories that fail to reach the goal state are excluded from metric computation to ensure that cost awareness is measured only on successful completions. To avoid being influenced by agents’ tool calling ability and reflect true cost-optimal planning capability, we also exclude failed tool calls. We thereby disentangle agents’ tool calling ability and cost-optimal planning ability.
    Task Completion metric.
    For Task Completion Ratio, we exclude all samples in which the agent fails to reach the goal state, since such predictions amount to mere guesses rather than answers derived from a complete tool-call trajectory.
@@
    B.3.3 Illustrative Examples
    To better understand the proposed metrics, we present a concrete example based on the trajectories shown in Figure 8. Let $\tau_{gt}$ denote the Ground-truth trajectory (green path), which employs a sequence of two composite tools with a total cost of $40.73+42.67=83.40$. In contrast, let $\tau_{agent}$ denote the Agent trajectory (red path), which adopts a less efficient plan consisting of three tools with a total cost of $22.22+38.55+23.56=84.33$. In the following, we illustrate our proposed metrics one by one using this example.
    (1) Cost Gap. This metric quantifies the economic inefficiency of the agent. In this scenario, the Cost Gap is calculated as the difference between the agent’s total cost and the ground-truth cost: 84.3383.40=0.9384.33-83.40=0.93. This positive value indicates that the agent failed to find the most cost-effective solution compared to $\tau_{gt}$.
    (2) Path Metrics (EMR, AED, ANED). These metrics assess the structural alignment between $\tau_{agent}$ and $\tau_{gt}$ . Specifically, the Exact Match Ratio (EMR) is 0 for this instance because the agent’s tool sequence ($\tau_{agent}$) differs markedly from the ground-truth sequence ($\tau_{gt}$). The Average Edit Distance (AED) measures the minimum number of edit operations required to transform $\tau_{agent}$ into $\tau_{gt}$. In this specific case, since the agent’s tool chain (Tool 3 $\to$ Tool 4 $\to$ Tool 5) is completely disjoint from the ground truth (Tool 1 $\to$ Tool 2), the transformation necessitates three operations. For instance, substituting Tool 3 with Tool 1, substituting Tool 4 with Tool 2, and deleting Tool 5 results in an edit distance of 3. The Average Normalized Edit Distance (ANED) normalizes the AED by the maximum length of the two trajectories (i.e., $\max(|\tau_{agent}|,|\tau_{gt}|)=\max(3,2)=3$). This yields an ANED of $3/3=1.0$, quantifying that the agent’s plan exhibits a 100% structural divergence from the optimal strategy. Note that the final reported values for these metrics are computed as the mean across all test instances.
    (3) Task Completion Ratio (TCR). Despite the sub-optimal path, the agent successfully reached the final goal state (TravelLocation). Therefore, the TCR for this sample is 1 (or 100%), highlighting that the agent was effective in task completion but lacked efficiency. The final reported TCR is calculated as the average over all samples.
    (4) Invalid Tool-Use Ratio (ITUR). This metric measures the proportion of failed tool invocations, such as calling non-existent tools, using wrong parameters, or violating prerequisites. In the example of Figure 8, all tool calls executed by the agent were valid transitions within the tool graph, resulting in an ITUR of 0. However, in cases where an agent attempts to invoke a tool without satisfying its input dependencies (e.g., calling a tool that requires LocationCandidate_L1 without first obtaining it), such calls would be counted as invalid failures.
@@ B.6.1 Deterministic Cost Assignment
    Here $a$ denotes an atomic tool, $T$ a composite tool with $k$ atomic components, and $h(\cdot)$ the hash-based seed derivation function. The $\sqrt{k}$ scaling in equation (2) reflects empirical observations that execution variance grows sublinearly with pipeline length Conway et al. (1967); Hopp and Spearman (2011).
@@ B.6.2 Deterministic Blocking Parameters
    Thus, this hierarchical seeding ensures that: (i) given $(S,q)$ and the blocking type, the entire blocking sequence and all associated parameters are fully determined (only the trigger time would be different, and we will explain its fairness in Appendix B.6.3); (ii) different queries exhibit different blocking patterns with the same $S$; and (iii) experimental results are perfectly reproducible, facilitating controlled comparisons across models and further analysis.
@@
    Repeated Calls $\downarrow$ Extra Calls $\downarrow$ Wrong Parameters $\downarrow$ Inaccessible Calls $\downarrow$
@@
    B.6.3 Evaluation Fairness
    Dynamic blocking mechanism ensures fairness across models.
    Among the four types of blocking events, all trigger times are determined based on the optimal path. Therefore, given identical cost assignments, the first trigger time remains fixed. However, the subsequent blocking time depends on the model’s behavior, because the state the model reaches during the first block determines the optimal path and thus the trigger time for the next blocking. We argue this is fair across models:
    • If a model follows the correct trajectory, both the ground truth and all agents are identically affected during the whole interaction trajectory, preserving fairness in evaluation.
    • Conversely, if a model already deviates from the optimal path, the model-dependent blocking trigger time has no impact, as the data point’s EM score would already be zero. Consequently, the EMR metric remains unaffected.
    • Moreover, as shown in Table 3, EMR and ANED/AED exhibit opposite trends, suggesting that the latter two effectively capture performance differences as well.
    Moreover, the blocking parameters are determined through a pseudo-random seeding mechanism (see details in Appendix B.6.2), making them deterministic and consistent across all models. Therefore, we consider our dynamic blocking mechanism to be fair across models.
    B.6.4 Evaluation Robustness
    Random seed mechanism promotes evaluation robustness.
    Building on the deterministic seeding framework introduced in Appendix B.6.2, our evaluation benefits from the diverse cost assignment and thereby achieves robustness. Specifically, the global seed $S$ is hierarchically propagated into per-query seeds $S_{q}$ and subsequently into tool-level seeds through the hash-based function $h(S,q,\text{name})$ (Eq. 1–3). This mechanism ensures that each query instantiates a distinct cost landscape: atomic tools receive independently sampled costs, and composite tools derive their perturbed costs via $\epsilon_{q,T}$, which yield sufficiently varied environments that reveal the agent’s true cost-awareness rather than memorization of any fixed structure. Additionally, we evaluate models under three independent runs (details in Appendix C.2), each initialized with a different global seed $S$. The consistency of results across these runs confirms that the benchmark remains robust to seed variation, while the hierarchical generation preserves controlled diversity across instances.
    Deterministic blocking content poses no memorization risk.
    Although the specific blocked tools and new cost values are pre-determined by a fixed seed for reproducibility, they remain completely invisible to the agent until runtime, and our evaluation architecture explicitly forbids any cross-instance memory or episodic recall. Each task is executed in full isolation with a fresh context, making pattern exploitation impossible.
    B.6.5 Discussions on Evaluation
    The necessity of introducing controlled uncertainty.
    While fully deterministic blocking (i.e., identical blocking parameters and trigger times) would benefit cross-model comparability, it would fundamentally limit the benchmark’s ability to evaluate adaptive planning. Real-world environments are inherently dynamic—cost structures shift, tools fail, and preferences change—requiring agents to replan under uncertainty rather than memorize fixed trajectories. If blocking events were static across all queries, a model could trivially overfit to CostBench through fine-tuning or pattern memorization, achieving superior scores without improving its robustness or generalization, as observed in prior benchmarks prone to leakage Wu et al. (2025b); Guo et al. (2025). Introducing controlled uncertainty addresses this issue: the content of each blocking event is deterministic under a global seed, ensuring reproducibility, while the timing adapts to the agent’s progress, preventing shortcut exploitation and preserving evaluation diversity. This hybrid design strikes a balance—maintaining fairness through seed-controlled reproducibility, while providing enough diversity to meaningfully assess an agent’s ability to detect disruptions, revise plans, and recover from unexpected environmental changes.
@@ Appendix C Additional Experiment Results
    C.1 Hardest Test on Static Settings
    In Figure 2, we analyze model performance in static settings, with task sequence lengths ranging from 5 to 8. We further report the detailed performance values in Table 9 for the most challenging case, task sequence = 8. We observe that under this setting, the EMR of all Qwen-series models drops to single digits. Gemini-2.5-Pro also exhibits a substantial degradation, falling from 83.73 at task sequence = 5 to 35.15, while only GPT-5 maintains an EMR of approximately 75%. These results indicate that current models struggle even with relatively simple forms of complexity scaling. Notably, we argue that comparisons among Qwen-series models in this setting are of limited practical significance, as their performance is nearly indistinguishable from a greedy policy, suggesting that their decisions are effectively equivalent to random guessing.
    C.2 Evaluation Robustness
    Models Radius (EMR) (%) Radius (ANED) (%) Radius (AED)
    Qwen3-8B 2.82 2.55 0.085
    Qwen3-14B 3.59 2.83 0.132
    Qwen3-32B 3.72 3.10 0.120
    Llama-3.1-8B-Instruct 2.02 2.07 0.112
    GLM-4.5 3.04 2.94 0.101
    Deepseek-V3.1 2.82 2.51 0.117
    Gemini-2.5-Pro 2.81 1.64 0.067
    Claude-Sonnet-4 3.86 2.96 0.105
    GPT-4o 2.55 2.23 0.110
    GPT-5 1.11 1.56 0.047
    Average 2.83 2.44 0.100
    Table 8: 95% Confidence Intervals for Main Metrics. We employ the Bootstrap method Justus et al. (2024) with 10,000 resampling iterations on the test set ($N=381$) to estimate statistical uncertainty. The reported Radius values denote the half-width of the 95% confidence interval (i.e., the $\pm$ value) for Exact Match Ratio (EMR), Average Normalized Edit Distance (ANED), and Average Edit Distance (AED).
    Theoretical Bounds.
    To ensure the statistical validity of our evaluation, we quantified the uncertainty associated with our metrics using the Bootstrap method Justus et al. (2024) with $10,000$ resampling iterations. As presented in Table 8, our measurement exhibits relatively high precision across all evaluated models. The 95% confidence intervals are remarkably tight, with an average half-width radius of only $2.83\%$ for Exact Match Ratio (EMR) and $2.44\%$ for Average Normalized Edit Distance (ANED). Furthermore, the radius for Average Edit Distance (AED) is negligible at $0.100$. These narrow bounds confirm that the observed performance differences between models are statistically significant and not artifacts of sampling noise, demonstrating the robustness of our benchmarking methodology.
    Empirical Results.
    To further empirically ensure the robustness of our conclusions, which is controlled by a random seed, we conducted three additional experiments using seeds 1000, 2000, and 3000 to evaluate model performance. As illustrated in Figure 9, the LLMs exhibit no significant performance variation across different seeds, with differences in ANED and EMR remaining within 5%, confirming low seed sensitivity.